# Taming Rectified Flow for Inversion and Editing

### v1

@@ 3.3 RF-Edit
Incorporating higher-order terms enables RF-Solver to significantly reduce errors in the ODE-solving process, improving both sampling quality and inversion accuracy. Furthermore, we extend the application of RF-Solver to real-world image and video editing tasks, which are more challenging than reconstruction. In these scenarios, maintaining the content and structure of the original image is crucial. For instance, when adding new objects to a source image, other unrelated regions should remain unaffected by the editing process. However, directly applying RF-Solver during the inversion and denoising stages may cause the model to be overly influenced by the target prompt, leading to unintended modifications in other parts of the image or video (e.g., altering the unrelated which is not mentioned in the editing prompt). This issue is common across various existing editing methods [45, 17, 49].
To address this problem, we propose RF-Edit, which incorporates features from the inversion process into the denoising procedure. Specifically, during the last $n$ steps of inversion, we extract and store the Value feature $\{\mathcal{{\widetilde{V}}}_{t_{k}}^{m}\}$ and $\{\mathcal{{\widetilde{V}}}_{t_{k}+\Delta t_{k}}^{m}\}$ from the self-attention layers in the last $M$ transformer blocks at each timestep $k$. Here, $k\in\{n,n+1,\cdots,N\}$ and $m\in\{1,\cdots,M\}$. This process can be formulated as follows:
@@
$\{\mathcal{{\widetilde{V}}}^{m}_{t_{k}+\Delta t_{k}}\}$ $=\text{Extract}\big{(}\boldsymbol{v}_{\theta}(\boldsymbol{\widetilde{Z}}_{t_{k}+\Delta t_{k}},t_{k}+\Delta t_{k})\big{)},$ (14)
where $\boldsymbol{\widetilde{Z}}_{t_{k}}$ and $\boldsymbol{\widetilde{Z}}_{t_{k}+\Delta t_{k}}$ denote the latents during inversion for RF-Solver.
During the first $n$ timesteps of denoising, considering the $m$th transformer block at the timestep $k$, the original self-attention in the network $\boldsymbol{v}_{\theta}$ can be formulated as:
@@
Figure 4: Qualitative Results of Image and Video Reconstruction. Our method (The second row) illustrates better performance compared with the rectified flow baselines (The third row) on both image and video reconstruction.
@@
The proposed RF-Edit framework enables high-quality editing while preserving structural information. Building on this concept, we design two sub-modules for RF-Edit, specifically tailored for image editing and video editing. For image editing, we use FLUX [1] as the backbone, which comprises several double blocks and single blocks. Double blocks independently modulate text and image features, while single blocks concatenate these features for unified modulation. In this architecture, RF-Edit shares features within the single blocks, as they capture information from both the source image and the source prompt, enhancing the ability of the model to preserve the structural information of the source image. For video editing, we employ OpenSora [2] as the backbone. The DiT blocks in OpenSora include spatial attention, temporal attention, and text cross-attention. Within this architecture, the structural information of the source video is captured in the spatial attention module, where we implement feature sharing.
@@ 4.1 Setup
Baselines. We select the vanilla Rectified Flow sampler as the primary baseline for all tasks. Additionally, for image editing, we compare our method with P2P [17], DiffEdit [12], SDEdit [33], PnP [49], Pix2pix [37] and RF-Inversion [45]. For video editing tasks, we compare our methods with FateZero [39], FLATTEN [10], COVE [51], RAVE [25], Tokenflow [15]. Detailed experimental settings of these methods are provided in the Appendix.
Figure 5: Qualitative Results of Text-to-Image Generation. With the RF-Solver, the model can generate images with higher quality (The first row) than baselines (The Second row).
Metric FID ($\downarrow$) Clip Score ($\uparrow$)
RF 26.55 31.01
Ours 25.45 31.09
Table 1: Quantitive results on Text-to-Image Generation. RF-Solver outperforms the vanilla RF sampler.
Figure 6: Qualititive Comparison of Image Editing. With RF-Solver and feature-sharing mechanism in RF-Edit, our method can successfully handle various kinds of image editing cases, outperforming the previous SOTA methods. Zoom-in for the best views.
Implementation Details. In the experiment, we adopt the guidance-distilled variant of FLUX [1] for image tasks and OpenSora [2] for video tasks. For sampling, we set the denoising step to 10 for our method. For reconstruction and editing, the inversion and denoising steps are set to 30. The derivative computation in RF-Solver requires an additional forward pass, resulting in the network needing to forward twice at each timestep. As a result, when comparing our method with the Rectified Flow baselines, we set the number of timestep for the vanilla Rectified Flow to be twice that of our method in order to maintain the same number of function evaluations (NFE) to ensure a fair comparison. Both the baseline and our method are conducted on a single A100 GPU with 40GB memory.
Evaluation Metrics For text-to-image sampling, we randomly select 10k images from the MSCOCO validation set [29] and report the FID [18] and Clip Score [40]. For the inversion and reconstruction task, we report the Mean Square Error (MSE), LPIPS [60], SSIM [54], and PSNR [24]. For image editing tasks, we report the Clip Score [40], which reflects whether the edited images align with the target prompt, and LPIPS [60], which reflects whether the edited images preserve the content of the source images. For video editing tasks, we adopt the metric proposed by [22], including Subject Consistency (SC), Motion Smoothness (MS), Aesthetic Quality (AQ), and Imaging Quality (IQ). SC and MS assess the temporal consistency of the edited video, while AQ and IQ assess the visual quality.
@@ 4.2 Text-to-image Sampling
We compare the performance of our method with the vanilla rectified flow on the text-to-image generation task. Both the quantitative (Tab. 1) and qualitative results (Fig. 5) demonstrate the superior performance of RF-Solver in fundamental T2I generation tasks, producing higher-quality images that align more closely with human cognition.
@@ 4.3 Inversion and Reconstruction
We conduct experiments on inversion and reconstruction for both image and video modalities, comparing our methods with the vanilla rectified flow sampler. For image inversion and reconstruction, we use images from the MSCOCO dataset, and for video inversion and reconstruction, we select videos from social media platforms such as TikTok and other publicly available sources [3, 4]. We use GPT-4.0 to annotate the content of the images and videos in detail, and then manually polish the GPT-generated content. These annotations are used as the source and target prompts for inversion and reconstruction.
Quantitive Comparison. The quantitative comparisons (Tab. 2) are conducted on both images and videos to illustrate the similarity between the source and reconstruction results. Our method demonstrates superior performance across all four metrics compared with vanilla rectified flow.
@@
Table 2: Quantitative Results on Inversion and Reconstruction. Our methods significantly improve the reconstruction accuracy of both images and videos.
Qualitative Comparison. RF-Solver effectively reduces the error in the solution of RF ODE, thereby increasing the accuracy of the reconstruction. As illustrated in Fig. 4(a), the image reconstruction results using vanilla rectified flow exhibit noticeable drift from the source image, with significant alterations to the appearance of subjects such as cake, church, train, and jeep. For video reconstruction, as shown in Fig. 4(b), the baseline reconstruction results suffer from distortion. In contrast, RF-Solver significantly alleviate these issues, achieving more satisfactory results.
Figure 7: Qualititive Comparison of Video Editing. The first video has 200 frames with a resolution of $512\times 512$, and the second video has 60 frames with a resolution of $1024\times 768$ (we compress the frames for a neat layout). Our method outperforms various previous SOTA video editing methods, especially excelling in dealing with complex prompts.
@@ 4.4 Editing
We conduct experiments to evaluate the image and video editing performance of our methods. Image editing usually involves replacing the subject in the image with another one, adding new items, and global editing. For the first two types of editing, the background of the source image is expected to remain unchanged after editing. For global editing such as style transfer, the overall structure of the source image is expected to remain unchanged. Compared to image editing, video editing is much more challenging due to the complexity of modeling temporal motions in videos. Recent mainstream video editing methods usually focus on replacing the subjects and performing global editing.
Quantitative Comparison. For image editing, we perform quantitative comparison between our methods and baselines (Tab. 3), reporting the Clip score and LPIPS. Our method outperforms all other methods in Clip score, indicating that the edited images align well with the user-provided prompts. For LPIPS, it is noted that PnP [49] has a much lower value than all other methods. Based on the qualitative results (Fig. 6), it can be seen that PnP is only suitable for editing cases that do not significantly modify the structure or shape of the source image (such as changing red roses into yellow sunflowers). For shape editing, such as modifying a car into a motorbike, PnP fails to edit this case, resulting in an image very similar to the source. Consequently, although PnP has the lowest LPIPS score, its clip score is the lowest.
@@
Table 3: Quantitative Results of Image Editing. RF-Edit can effectively edit the images according to the prompts while keeping the unrelated regions unchanged.
For video editing, we compare our methods with baseline methods using the VBench [22] metrics (Tab. 4). The results illustrate that our methods successfully maintain temporal consistency in long videos, achieving the highest Subject Consistency (SC) and Motion Smoothness (MS) scores. Additionally, our method demonstrates superior visual quality, outperforming the baselines at Aesthetic Quality (AQ) and Image Quality (IQ).
@@
Table 4: Quantitative Results of Video Editing. RF-Edit outperforms a number of previous SOTA video editing methods.
Qualitative Comparison. For image editing, we compare the performance of our methods with several baselines across different types of editing tasks (Fig. 6). The baseline methods often suffer from background changes or fail to perform the desired edits. In contrast, our methods demonstrate satisfying performance, effectively achieves a balanced trade-off between the fidelity to the target prompt and preservation of the source image. To be noticed, although RF-inversion [45] also uses the rectified flow model for image editing (third row in Fig. 6), the structure of the source image which is unrelated to editing prompt (such as background and human appearance) is modified obviously.
For video editing, we primarily evaluate the performance of our methods on long videos (200 frames) and high-resolution videos ($1280\times 768$). Furthermore, we assess the performance on complicated videos and prompts where there are multiple objects in the video, and the user has different editing requirements for each object. The qualitative results are shown in Fig. 7. Our method successfully handles complicated editing cases (e.g., modifying the leftmost lion among three lions into a white polar bear and changing the other two small lions into orange tiger cubs), whereas all other baseline methods fail in this scenario. Our method also demonstrates strong performance in global editing tasks, such as transforming scenes into autumn.
4.5 Ablation Study
Taylor Expansion Order of RF-Solver. We investigated the impact of the Taylor expansion order in RF-Solver (Tab. 5) under the same NFE across different orders. The second-order expansion demonstrated a significant improvement across various tasks compared to the first-order expansion (i.e., the vanilla rectified flow). However, higher-order expansions did not yield further enhancements. We speculate that this is primarily due to higher-order Taylor expansions requiring more inference steps per timestep. With a fixed NFE, this results in a reduced overall number of timesteps compared to lower-order expansions, leading to suboptimal performance. Moreover, computing the higher-order derivatives of $v_{\theta}{t_{i}}$ substantially increases the complexity of the algorithm, posing challenges for practical applications. Consequently, we predominantly employed second-order expansion in our experiments.
@@ 4.5 Ablation Study
Table 5: Ablation Study on the Taylor Expansion Order. We choose the 2-order expansion (i.e. RF-Solver-2) for the downstream tasks for its effectiveness and simplicity.
Figure 8: Ablation Study of Feature-Sharing Step in RF-Edit. A too-small feature-sharing step leads to inconsistency between source and target images. On the other hand, a too-large feature-sharing step can cause the failure of editing.
@@ 5 Conclusion
In this paper, we propose RF-Solver, a versatile sampler for the rectified flow model that solves the rectified flow ODE with reduced error, thus enhancing the image and video generation quality across various tasks such as sampling and reconstruction. Based on RF-Solver, we further propose RF-Edit, which achieves high-quality editing performance while effectively preserving the structural information in source images or videos. Extensive experiments demonstrate the versatility and effectiveness of our methods.

### v2

@@ 3.3 RF-Edit
Incorporating higher-order terms enables RF-Solver to significantly reduce errors in the ODE-solving process, thereby enhancing both sampling quality and inversion accuracy. Furthermore, we extend the application of RF-Solver to the more complex real-world image and video editing tasks, which present greater challenges than reconstruction. In such scenarios, preserving the content and structure of the original image is crucial. For example, when replacing an object in a source image with another one, regions unrelated to the object in this image are expected to remain unaffected by the editing process. However, directly applying RF-Solver during the inversion and denoising stages may cause the model to be overly influenced by the target prompt, resulting in unintended modifications in other regions of the source image or video. Similar issues are common across various existing editing methods [58, 20, 64].
To address this problem, we propose RF-Edit, which builds upon the diffusion transformer architecture. Specifically, we focus on the self-attention layer in the last $M$ transformer blocks of $\boldsymbol{v}_{\theta}$ at the last $n$ timesteps during inversion. The self-attention operation can be formulated by:
$\boldsymbol{\widetilde{F}}^{m}_{t_{k}}=\text{Attention}(\mathcal{{\widetilde{Q}}}_{t_{k}}^{m},\mathcal{{\widetilde{K}}}_{t_{k}}^{m},\mathcal{{\widetilde{V}}}_{t_{k}}^{m}).$ (10)
Here, $k\in\{N-n,\cdots,N\}$, and $m\in\{1,\cdots,M\}$, $\boldsymbol{\widetilde{F}}^{m}_{t_{k}}$ denotes the output feature of the self-attention module and $\mathcal{{\widetilde{Q}}}_{t_{k}}^{m},\mathcal{{\widetilde{K}}}_{t_{k}}^{m},\mathcal{{\widetilde{V}}}_{t_{k}}^{m}$ represent query, key and value for attention during the inversion process, respectively. We extract and store the Value feature $\{\mathcal{{\widetilde{V}}}_{t_{k}}^{m}\}$ and $\{\mathcal{{\widetilde{V}}}_{t_{k}+\Delta t_{k}}^{m}\}$ in the process of RF-Solver algorithm (Algorithm 1):
@@
$\{\mathcal{{\widetilde{V}}}^{m}_{t_{k}+\Delta t_{k}}\}$ $=\text{Extract}\big{(}\boldsymbol{v}_{\theta}(\boldsymbol{\widetilde{Z}}_{t_{k}+\Delta t_{k}},t_{k}+\Delta t_{k})\big{)}.$ (12)
During the first $n$ timesteps of denoising, considering the $m$th transformer block at the timestep $k$, the original self-attention can be formulated as:
@@
Figure 4: Qualitative results of image and video reconstruction. Our method (the second row) demonstrates superior performance compared to the vanilla rectified flow baselines (the third row).
@@
The proposed RF-Edit framework enables high-quality editing while effectively preserving the structural information of the source image/video. Building on this concept, we design two sub-modules for RF-Edit, specifically tailored for image editing and video editing (Fig. 3). For image editing, RF-Edit employs FLUX [2] as the backbone, which comprises several double blocks and single blocks. Double blocks independently modulate text and image features, while single blocks concatenate these features for unified modulation. In this architecture, RF-Edit shares features within the single blocks, as they capture information from both the source image and the source prompt, enhancing the ability of the model to preserve the structural information of the source image. For video editing, RF-Edit employs OpenSora [3] as the backbone. The DiT blocks in OpenSora include spatial attention, temporal attention, and text cross-attention. Within this architecture, the structural information of the source video is captured in the spatial attention module, where we implement feature sharing.
@@ 4.1 Setup
Baselines. We use the vanilla Rectified Flow sampler (Euler Sampler) as the primary baseline for all tasks. Besides, for sampling and inversion, we compare our method with DPM-Solver++ [44] and the Heun sampler (a 2-order ODE solver). For image editing, we compare our method with P2P [20], DiffEdit [11], SDEdit [45], PnP [64], Pix2pix [50] and RF-Inversion [58]. For video editing tasks, we compare our method with FateZero [53], FLATTEN [9], COVE [68], RAVE [30], Tokenflow [16]. Detailed experimental settings of baselines are provided in the Appendix.
Figure 5: Qualitative results of text-to-image generation. By employing the RF-Solver, the model is able to generate images with higher quality (the first row) than baselines (the second row).
Implementation Details. In the experiment, we adopt the guidance-distilled variant of FLUX [2] for image tasks and OpenSora [3] for video tasks. The derivative computation in RF-Solver requires an additional forward pass, resulting in the network needing to forward twice at each timestep. As a result, when comparing our method with the Rectified Flow baselines, we set the number of timesteps for the vanilla Rectified Flow to be twice that of our method to ensure a fair comparison under the same number of function evaluations (NFE). More information is provided in the Appendix.
Evaluation Metrics For text-to-image sampling, we randomly select 10k images from the MSCOCO validation set [39] and report the FID [21] and CLIP Score [54]. For the inversion and reconstruction task, we report the Mean Square Error (MSE), LPIPS [80], SSIM [72], and PSNR [28]. For image editing tasks, we report the CLIP Score [54] and LPIPS [80]. For video editing tasks, we adopt the metric proposed by [26], including Subject Consistency (SC), Motion Smoothness (MS), Aesthetic Quality (AQ), and Imaging Quality (IQ). Detailed explanations of these metrics are provided in the Appendix.
@@ 4.2 Text-to-Image Sampling
We compare the performance of our method with DPMSolver++ [44], the vanilla RF sampler, and Heun sampler on the text-to-image generation task. Both the quantitative (Tab. 1) and qualitative results (Fig. 5) demonstrate the superior performance of RF-Solver in fundamental T2I generation tasks, producing higher-quality images that align more closely with human cognition.
Figure 6: Qualitative comparison of image editing. With RF-Solver and feature-sharing mechanism in RF-Edit, our method can successfully handle various kinds of image editing cases, outperforming the previous SOTA methods. Zoom in for the best views.
DPMSolver++ RF RF-Heun Ours
FID ($\downarrow$) 24.63 25.33 24.40 24.03
CLIP Score ($\uparrow$) 30.62 31.01 31.03 31.09
Table 1: Quantitative results on text-to-image generation. RF-Solver outperforms several baselines.
@@ 4.3 Inversion and Reconstruction
We conduct experiments on inversion and reconstruction for both image and video modalities, comparing our method with the vanilla RF sampler and the Heun sampler.
Quantitative Comparison. The quantitative comparisons (Tab. 2) are conducted to reflect the similarity between the source and reconstruction results. Our method demonstrates superior performance across all four metrics compared with the vanilla RF sampler and Heun sampler.
@@
Table 2: Quantitative results on inversion and reconstruction. Our method significantly improves the accuracy of reconstruction for both images and videos.
Qualitative Comparison. RF-Solver effectively reduces the error in the solution of RF ODE, thereby increasing the accuracy of the reconstruction. As illustrated in Fig. 4(a), the image reconstruction results using vanilla rectified flow exhibit noticeable drift from the source image, with significant alterations to the appearance of subjects in the image. For video reconstruction, as shown in Fig. 4(b), the baseline reconstruction results suffer from distortion. In contrast, RF-Solver significantly alleviates these issues, achieving more satisfactory results.
Figure 7: Qualitative comparison of video editing. The first video comprises 200 frames with a resolution of $512\times 512$, while the second video contains 60 frames with a resolution of $1024\times 768$ (frames are compressed for a neat layout in the figure).
@@ 4.4 Editing
We conduct experiments to evaluate the image and video editing performance of our method. Image editing usually involves replacing the subject in the image with another one, adding new items, and global editing. For the first two types of editing, the background of the source image is expected to remain unchanged after editing. For global editing such as style transfer, the overall structure of the source image is expected to remain unchanged. Recent mainstream video editing methods usually focus on replacing the subjects and performing global editing for the source video.
Quantitative Comparison. In image editing, Our method outperforms all other methods in CLIP score (Tab. 3), indicating that the edited images align well with the user-provided prompts. For LPIPS, it is noted that PnP [64] has a much lower value than all other methods. Based on the qualitative results (Fig. 6), it can be seen that PnP is only suitable for editing cases that do not significantly modify the structure or shape of the source image (such as changing red roses into yellow sunflowers). It fails in the case of shape editing, resulting in an image very similar to the source. Consequently, although PnP has the lowest LPIPS score, its CLIP score is the lowest.
@@
Table 3: Quantitative results of image editing. RF-Edit effectively edit the images according to the prompts while preserving the integrity of unrelated regions.
For video editing, RF-Edit achieves higher scores on the popular VBench [26] metrics (Tab. 4). The results illustrate that our method successfully maintains temporal consistency while demonstrating superior visual quality.
@@
Table 4: Quantitative results of video editing. RF-Edit outperforms several previous SOTA video editing methods.
Qualitative Comparison. For image editing, we compare the performance of our method with several baselines across different types of editing tasks (Fig. 6). The baseline methods often suffer from background changes or fail to perform the desired edits. In contrast, our method demonstrates satisfying performance, effectively achieving a balanced trade-off between the fidelity to the target prompt and the preservation of the source image.
For video editing, we primarily evaluate the performance of our method on long videos (200 frames) and high-resolution videos ($1280\times 768$). The qualitative results are shown in Fig. 7. RF-Edit illustrates impressive performance in handling complicated editing cases (e.g., modifying the leftmost lion among three lions into a white polar bear and changing the other two small lions into orange tiger cubs), whereas all other baseline methods fail in this scenario. RF-Edit also demonstrates strong performance in global editing tasks, such as transforming scenes into autumn.
@@
Table 5: Ablation study on the Taylor Expansion order. We select the 2-order expansion (i.e. RF-Solver-2) for various downstream tasks due to its effectiveness and simplicity.
4.5 Ablation Study
Taylor Expansion Order of RF-Solver. We investigated the impact of the Taylor expansion order in RF-Solver (Tab. 5) under the same NFE across different orders. The second-order expansion demonstrated a significant improvement across various tasks compared to the first-order expansion (i.e., the vanilla rectified flow). However, higher-order expansions do not yield further enhancements. We speculate that this is primarily due to higher-order Taylor expansions requiring more inference steps per timestep. With a fixed NFE, this results in a reduced overall number of timesteps compared to lower-order expansions, leading to suboptimal performance. Moreover, computing the higher-order derivatives of $\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t_{i}},{t_{i}})$ substantially increases the complexity of the algorithm, posing challenges for practical applications. Consequently, we predominantly employed second-order expansion in our experiments.
Figure 8: Ablation study of feature-sharing step in RF-Edit. A too-small feature-sharing step results in the inconsistency between source and target images. Conversely, a too-large feature-sharing step can lead to the failure of editing.
@@ 5 Conclusion
In this paper, we propose RF-Solver, a versatile sampler for the rectified flow model that solves the rectified flow ODE with reduced error, thus enhancing the image and video generation quality across various tasks such as sampling and reconstruction. Based on RF-Solver, we further propose RF-Edit, which achieves high-quality editing performance while effectively preserving the structural information in source images or videos. Extensive experiments demonstrate the versatility and effectiveness of our method.
Supplementary Material
6 Experimental Settings
6.1 Baselines and Implementation Details
Text-to-Image Generation. We compare our methods with the following baselines: FLUX with the vanilla sampler, Heun Solver, and DPM-Solver. The Heun Solver is a second-order ODE solver that can be applied to pretrained rectified flow to solve the ODE more precisely. DPM-Solver is a high-order sampler for diffusion ODE, which is not suitable for RF-based models like FLUX. As an alternative, we apply the DPM-Solver on Stable Diffusion to evaluate its performance. For FLUX with the vanilla sampler and the Heun Solver, we randomly select 10000 images from the MS-COCO validation dataset and use their caption as the prompt for generation. The resolution of generated images is $1024\times 1024$. For DPM-Solver, we adopt the implementation from the diffuser, adopting its default setting to generate images. The total NFE for generating one image is set to 10 for both our method and baselines.
Inversion. We compare the performance of our methods among RF with the vanilla sampler and the Heun sampler. For image inversion, we also use the images from the MS-COCO validation set. For video inversion, we select videos from social media platforms such as TikTok and other publicly available sources. We have observed the quality of the text prompts significantly influence the quality of inversion. Consequently, we employ GPT-4o to generate detailed captions for both images and videos, which are then used in the inversion tasks. The total NFE for generating one image/video is set to 50 for both our method and baselines.
Editing. For image editing, we compare our methods with RF-inversion and several diffusion-based editing methods. For RF-inversion, we adopt the implementation in ComfyUI [1]. For other baselines, we use their implementation from diffusers. For each baseline, we adjust the relevant hyper-parameters to achieve optimal results. For video editing, we use the official codes of all the baseline methods and tune the hyper-parameters to achieve satisfactory results. For image editing, we share the features of the last 19 single blocks in FLUX. For video editing, we share the features of the last 14 blocks in Open-Sora. We adjust the hyper-parameter of feature-sharing steps to achieve better results for both image and video editing.
Figure 9: Stylization Results.
6.2 Evaluation Metrics
For text-to-image sampling, we report Fréchet Inception Distance (FID) and CLIP Scores. The FID is a metric used to evaluate the quality of generated images by assessing the similarity between the distributions of real and generated image features, typically extracted using a pre-trained Inception network. The CLIP Score evaluates the alignment between generated images and textual descriptions by measuring the similarity of their embeddings within a shared multimodal space using the CLIP model.
For Inversion tasks, our evaluation metrics include MSE, LPIPS, SSIM, and PSNR. MSE measures the average squared difference between predicted and ground-truth values, quantifying the overall error in pixel intensity. LPIPS assesses perceptual similarity between images by comparing deep feature representations extracted from neural networks, aligning with human perception. SSIM evaluates image quality by comparing luminance, contrast, and structure to measure the similarity between the reference and reconstructed images. PSNR quantifies the ratio between the maximum possible signal value and the power of noise, commonly used to assess image reconstruction quality.
For video editing, we adopt the VBench Metrics [26]. The evaluation criteria include Subject Consistency, Motion Smoothness, Aesthetic Quality, and Imaging Quality. Subject Consistency measures whether the subject (e.g., a person) remains consistent throughout the video by computing the similarity of DINO features [6] across frames. Motion Smoothness assesses the smoothness of motion in the generated video using motion priors from the video frame interpolation model [38]. Aesthetic Quality evaluates the artistic and visual appeal of each frame as perceived by humans, leveraging the LAION aesthetic predictor [34]. Imaging Quality examines the level of distortion in the generated frames (e.g., blurring or flickering) based on the MUSIQ image quality predictor [32].
7 Stylization
We provide more results on image stylization (Fig. 9). In this task, we perform inversion on the image with a reference style, following the same overall pipeline as in image editing. RF-Edit demonstrates superior performance in stylized generation, effectively preserving the style of the source image while generating images that align with the user-provided prompts.
8 Limitation
Despite achieving impressive performance across various tasks, our methods occasionally exhibit performance instability, particularly in video editing. Additionally, the FLUX and OpenSora models contain a large number of parameters, leading to significant memory consumption, which may limit the applicability of our method on resource-constrained devices. To mitigate these issues, we plan to further optimize our method and explore its efficacy on more lightweight models.